OpenStudy (kaylala):

Find the Exact Value of: tan 7π/12 by using the values of 3π/4 and 2π/3

OpenStudy (kaylala):

@phi

OpenStudy (kaylala):

@Mertsj

OpenStudy (kaylala):

@Loser66 @jagatuba @e.mccormick @eliassaab

OpenStudy (kaylala):

@Compassionate @inkyvoyd @nincompoop @Preetha @wio

OpenStudy (anonymous):

is it tan 17 pi/12 or tan 7 pi/12 ?

OpenStudy (kaylala):

@surjithayer ^

OpenStudy (kaylala):

@jim_thompson5910

OpenStudy (kaylala):

@tkhunny

OpenStudy (kaylala):

@shamil98 @timo86m @ranga @charlotte123 @beccaboo333

OpenStudy (beccaboo333):

No need to tag so many people. I'm sorry I'm unable to help.

OpenStudy (mertsj):

\[\tan (A+B)=\frac{\tan A+\tan B}{1-\tan A \tan B}\]

OpenStudy (mertsj):

So : \[\tan \frac{17\pi}{12}=\tan (\frac{3\pi}{4}+\frac{2\pi}{3})\]

OpenStudy (mertsj):

Plug into the formula.

OpenStudy (kaylala):

pardon, but i dont know what's next. i'm new to this @Mertsj

OpenStudy (mertsj):

Did you see the formula I posted?

OpenStudy (mertsj):

Replace each A with: \[\frac{3\pi}{4}\] Replace each B with:\[\frac{2\pi}{3}\]

OpenStudy (kaylala):

yes i did see it

OpenStudy (kaylala):

okay. let me try

OpenStudy (kaylala):

what's next after plugging the values in?? i already did it @Mertsj

OpenStudy (mertsj):

What is the tangent of 3pi/4?

OpenStudy (kaylala):

none? it's not in the graph

OpenStudy (mertsj):

What is the tangent of pi/4?

OpenStudy (mertsj):